Required information

The Foundational 15 [LO13-1, LO13-2, LO13-3, LO13-5, LO13-6]

[The following information applies to the questions displayed below.]

Cardinal Company is considering a five-year project that would require a $2,855,000 investment in equipment with a useful life of five years and no salvage value. The company’s discount rate is 14%. The project would provide net operating income in each of five years as follows:

Sales $ 2,867,000
Variable expenses 1,125,000
Contribution margin 1,742,000
Fixed expenses:
Advertising, salaries, and other fixed out-of-pocket costs $ 706,000
Depreciation 571,000
Total fixed expenses 1,277,000
Net operating income $ 465,000

5. What is the project profitability index for this project? (Round your answer to 2 decimal places.)
Answer #1

Annual Cash inflow = 465000+571000 = 1036000

Profitability index = Present value of cash inflow/Present value of cash outflow

                           = (1036000*3.433)/2855000

Profitability index = 1.25:1